OpenStudy (anonymous):

A body P has KE E . Another body Q whose mass is 9 times than P also has a KE E . Calculate the ratio of velocities of P and Q ?

OpenStudy (anonymous):

@Yahoo!

OpenStudy (anonymous):

Let the Mass of body p = x mass of q = 9x\[E = \frac{ 1 }{2 } * x * v1^2\] \[E = 1/2 * 9x * v2^2\]

OpenStudy (anonymous):

\[\frac{ v1^2 }{ v2^2 } = \frac{ 2E }{ X } * \frac{ 9X }{ 2E }\]

OpenStudy (anonymous):

Nw.....Just Solve..)

OpenStudy (anonymous):

Did u understand @ashna

OpenStudy (anonymous):

not the comment above " now just solve "

OpenStudy (anonymous):

Did u understand???

OpenStudy (anonymous):

you have asked me to solve , i dont know how to do that ?

OpenStudy (anonymous):

\[\frac{ v1^2 }{ v2^2 } = \frac{ 9 }{ 1 }\] \[\frac{ v1 }{ v2 } = 3\]

OpenStudy (anonymous):

is it over ?

OpenStudy (anonymous):

yup.....The ratio is 3 : 1

OpenStudy (anonymous):

yeah .. thanks :)

OpenStudy (anonymous):

Welcome...